ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 39 Werbowski, Jarosław, Poznań University of Technology, Poland Oscillatory solutions of functional equations The functional equation

Q_{0}(t)x(t) + Q_{1}(t)x(g(t)) + … + Q_{m+1}x(g^{m+1}(t)) = 0, m ≥ 1,

where Q_{i} : I → R (i = 0, 1, …, m + 1), I ⊂ R_{+} = [0, ∞) and g : I → I are given functions is considered. Sufficient conditions of this equation are given. Part of these results is joint work with W. Nowakowska.